OpenStudy (anonymous):

The dimensions of a smaller rectangle are 3 ft by 9 ft. The dimensions of a larger rectangle are 5 ft by 11 ft. Find the ratio of the perimeter of the larger rectangle to the perimeter of the smaller rectangle

OpenStudy (anonymous):

Perimeter of a rectangle is just adding up the side lengths all the way around. For the first rectangle, so perimeter of small = 3+3+9+9 = 22 feet For the second rectangle, perimeter of large = 5+5+11+11 = 32 feet A ratio of small:large is 22:32.